Please help me solve this problem:
0.25x+.05(8-x)=.10(-28)

Answer by checkley77(12844)

0.25x+.05(8-x)=.10(-28)
.25X+.4-.05X=-2.8
.2X=-2.8-.4
.2X=-3.2
X=-3.2/.2
X=-16 ANS.
PROOF:
.25*-16+.05(8+16)=-2.8
-4+.05*24=-2.8
-4+1.2=-2.8
-2.8=-2.8